# Fast post-process Bayesian inference with Variational Sparse Bayesian Quadrature

### v2

@@ C.6 Runtime analysis
C.7 Posterior estimation: MCMC or variational inference?
After having fitted a surrogate regression model (a sparse GP or a neural network) to the log-density function, the key issue is how to obtain an estimate of the posterior density (see Figure 1 in the main text). In the paper, we demonstrated how (stochastic) variational inference on the surrogate (svi) can successfully recover the posterior, based on prior work [Acerbi, 2018]. However, why not directly run mcmc on the log-density surrogate to obtain approximate posterior samples, as done by other works [Rasmussen, 2003, Nemeth and Sherlock, 2018, Järvenpää et al., 2021]?
The answer is that we attempted to use mcmc (slice sampling; Neal, 2003) and empirically found that, for both sparse GPs and neural networks, mcmc often gives inferior results compared to svi. For example, Table S4 shows the performance metrics for sgpr with svi (i.e., vsbq), sgpr with mcmc, nnr with svi, and nnr with mcmc, on the Bayesian timing model benchmark problem. mcmc is substantially less robust and yields worse results compared with svi.
@@ C.7 Posterior estimation: MCMC or variational inference?
Instead, for variational inference, initializing the variational distribution components around the high-density observation points and imposing an additional penalty loss in the elbo optimization for bounding the variational parameters (means and scales of the mixture components), as done in Acerbi [2020], help constrain the variational distribution in the local trust region. As a further advantage, for a sparse GP, variational inference with Bayesian quadrature is efficient and affords computation of the elbo and its standard deviation, $\textsc{elbo}_{\text{sd}}$, for assessing its uncertainty.
@@
C.8 Note on package versions

### v3

@@ C.7 Runtime analysis
C.8 Posterior estimation: MCMC or variational inference?
After having fitted a surrogate regression model (a sparse GP or a neural network) to the log-density function, the key issue is how to obtain an estimate of the posterior density (see Figure 1 in the main text). In the paper, we demonstrated how (stochastic) variational inference (svi) on the surrogate can successfully recover the posterior, based on prior work [Acerbi, 2018]. [^5] In this paper, svi refers to stochastic variational inference with the reparameterization trick, including advi. However, why not directly run mcmc on the log-density surrogate to obtain approximate posterior samples, as done by other works [Rasmussen, 2003, Nemeth and Sherlock, 2018, Järvenpää et al., 2021]?
The answer is that we attempted to use mcmc (slice sampling; Neal, 2003) and empirically found that, for both sparse GPs and neural networks, mcmc often gives inferior results compared to svi. For example, Table S5 shows the performance metrics for sgpr with svi (i.e., vsbq), sgpr with mcmc, nnr with svi, and nnr with mcmc, on the Bayesian timing model benchmark problem. mcmc is substantially less robust and yields worse results than svi.
@@ C.8 Posterior estimation: MCMC or variational inference?
Instead, for variational inference, initializing the variational distribution components around the high-density observation points and imposing an additional penalty loss in the elbo optimization for bounding the variational parameters (means and scales of the mixture components), help constrain the variational distribution in the local trust region, see Section C.1 for details. As a further advantage, for a sparse GP, variational inference with Bayesian quadrature is efficient and affords computation of the elbo and its standard deviation, $\textsc{elbo}_{\text{sd}}$, for assessing its uncertainty.
@@
C.9 Note on package versions